y(1/2) = -2, Solve the initial value problem: 9y + 18y + 14y = 0, y (1/2) = -1. Give your answer as y=... . Use x as the i
Answer #1

Y CT/2) = -2 qy +184 +144 = 0 ya Mra) = -1 (90? +180 +14)y=o Auxulary eqution . qm2+18m+ 14 =o ma- -18 + √(1812-4 (147691 2from of ③ © terases para ) es uairpom = (-11-E-le- (0.1-8-742 Cu firmus (15-7-2423+ (v.)m Mestlung -) ?? +(1.6) (e)> (45-15-I c): Tea) (9-25) Cos (px) + elected (52) ( 24 25 s24.com enten (se (5)(938) 09( 462) + Came( 53*) (*son qox)] * yo eh (se($